### v1

# X-Dancer: Expressive Music to Human Dance Video Generation

Abstract
We present X-Dancer, a novel zero-shot music-driven image animation pipeline that creates diverse and long-range lifelike human dance videos from a single static image. As its core, we introduce a unified transformer-diffusion framework, featuring an autoregressive transformer model that synthesize extended and music-synchronized token sequences for 2D body, head and hands poses, which then guide a diffusion model to produce coherent and realistic dance video frames. Unlike traditional methods that primarily generate human motion in 3D, X-Dancer addresses data limitations and enhances scalability by modeling a wide spectrum of 2D dance motions, capturing their nuanced alignment with musical beats through readily available monocular videos. To achieve this, we first build a spatially compositional token representation from 2D human pose labels associated with keypoint confidences, encoding both large articulated body movements (e.g., upper and lower body) and fine-grained motions (e.g., head and hands). We then design a music-to-motion transformer model that autoregressively generates music-aligned dance pose token sequences, incorporating global attention to both musical style and prior motion context. Finally we leverage a diffusion backbone to animate the reference image with these synthesized pose tokens through AdaIN, forming a fully differentiable end-to-end framework. Experimental results demonstrate that X-Dancer is able to produce both diverse and characterized dance videos, substantially outperforming state-of-the-art methods in term of diversity, expressiveness and realism.
{strip}[Uncaptioned image]
Figure 1: We present X-Dancer, a unified transformer-diffusion framework for zero-shot, music-driven human image animation from a single static image, capable of handling diverse body forms and appearances. Our method enables the synthesis of highly expressive and diverse full-body dance motions that are synchronized with music, including detailed movements at the head and hands, which are then seamlessly translated into vivid and lifelike dance videos. Code and model will be available for research purposes.
1 Introduction
Dance is a universal form of self-expression and social communication, deeply embedded in human behavior and culture. With the rise of social media platforms like TikTok and YouTube Shorts, people increasingly share self-expressive dance videos online. However, creating expressive choreography typically demands practice and even professional training. From a computational aspect, generating realistic dance movements is challenging due to the freeform, personalized nature of dance and its alignment with musical rhythm and structure. In this work, we tackle the challenge of creating continuous, expressive and lifelike dance videos from a single static image, driven solely by a music track.
This study addresses two key challenges in music-driven human image animation: (1) generating smooth, diverse full-body movements at finer scales that capture the complex, non-linear synchronization with music inputs, and (2) translating these generated body movements into high-fidelity video outputs that maintain visual consistency with the reference image and ensure temporal smoothness. Prior approaches [32, 37] mainly focus on computationally generating 3D human poses, such as SMPL skeletons [26], from music inputs, utilizing diffusion- or GPT-based frameworks. While these methods excel in producing high-quality, clean motions, they are constrained by limited training datasets (primarily the multi-view AIST dataset [38, 21]), which lack diversity, and contain 3D body poses only (excluding head and hands motions). Expanding these datasets with widely available 2D monocular dance videos would require 3D human pose estimation, which is often error-prone and risks degraded motion quality and consistency. Moreover, we target at photorealistic dance video generation rather than 3D skeleton or mesh animations.
With recent advances in diffusion models that have achieved unprecedented diversity and quality in image and video generation, numerous works have leveraged their generative capabilities to synthesize visually compelling videos by animating a reference image with motion signals such as 2D skeletons [15, 6], dense poses [43], and depth maps [52]. Unlike motion transfer settings that derive motion from a driving video, our goal is to generate motion signals that are consistent with the reference body shape and aligned with the input musical beats. Recently, a few studies [42, 35] have attempted to synthesize visual outputs end-to-end from audio inputs using diffusion models. While these methods have advanced in realism and dynamic quality, they still struggle to capture long-range motion and audio context due to high computational demands. Moreover, these frameworks have primarily been validated on audio-driven portrait animations, leaving uncertain whether they can accommodate the complexities of full-body kinematics and rapid motion transitions.
To this end, we propose X-Dancer, a unified framework that integrates an autoregressive transformer model for generating extended dance sequences attuned to input music, coupled with a diffusion model to produce high-resolution, lifelike videos. In contrast to prior methods focused on 3D music-to-dance motion generation, our approach models 2D human motion, leveraging widely accessible dance videos where 2D pose estimation is both reliable and scalable. For effective autoregressive motion generation, we develop a multi-part tokenization scheme for per-frame 2D whole-body poses, incorporating detected keypoint confidences to capture multi-scale human motions with motion blur and various occlusions. Thereafter we train a cross-modal transformer that auto-regressively predicts future N-frame pose tokens, paired with per-frame music features extracted with Jukebox [9] and Librosa [19]. Our design enables the model to capture a broader diversity of expressive, music-aligned movements with enhanced scalability in both model complexity and data scale. We then leverage a T2I diffusion model with temporal layers to animate the reference image by implicitly translating the generated confidence-aware motion tokens into spatial guidance via a trainable motion token decoder. Specifically, it upsamples from a learnable feature map into multi-scale spatial feature guidance, integrating motion tokens through AdaIN [16] during the upsampling process. By co-learning pose translation with temporal and appearance reference modules [12, 5], the diffusion backbone interprets motion tokens with temporal context and body shape reference, leading to better shape-disentangled pose control, smoother and more robust visual outputs even under low-confidence or jittering pose sequences. This design establishes an end-to-end transformer-diffusion learning framework, merging the transformer’s strengths in long-context understanding and generation with the diffusion model’s high-quality visual synthesis.
To the best of our knowledge, X-Dancer represents the first music-driven human image animation framework. Trained on a curated dataset of 100K music-dance video clips, our method excels at generating diverse, expressive and detailed whole-body dance video animations attuned to input music, adaptable to both realistic and stylized human images across various body types. We comprehensively evaluate our model on challenging benchmarks, and X-Dancer outperforms state-of-the-art baselines both quantitatively and qualitatively. Additionally, we highlight its scalability and customization capabilities, showcasing scalability across different model and data scales, as well as fine-tuning to specific choreography. We summarize our contributions as follows,
  • A novel transformer-diffusion based music-to-dance human image animation, achieving state-of-the-art performance in terms of motion diversity, expressiveness, music alignment and video quality.
  • A cross-modal transformer model that captures long-range synchronized dance movements with music features, employing a multi-scale tokenization scheme of whole-body 2D human poses with keypoint confidence.
  • A diffusion-based human image animation model that interprets temporal pose tokens and translates them into consistent high-resolution video outputs.
  • Demonstration of captivating zero-shot music-driven human image animations, along with characterized choreography generation.
2 Related Work
2.1 Music to Dance Generation
Significant progress has been made in realistic human motion generation [51, 32, 37, 27, 7, 48] from various inputs, such as input speech [11, 30, 25, 47, 2] or text [34, 17, 29, 50]. However, the task of music-to-dance generation [32, 1, 37, 24, 23, 20, 28, 46] presents unique challenges: (1) ensuring the generated dance rhythmically aligns with the music, and (2) producing intricate motions with diverse styles and speeds. Due to data limitations, several 3D human pose datasets [21, 22, 20, 28] have been proposed. The AIST++ dataset [21] is a 3D music pose dataset containing 1,408 dance motions tailored to various music genres. FineDance [22], a 3D dataset focusing on fine-grained hand motions, includes 14 hours of dance data. Models like Bailando [32] and EDGE [37] have leveraged AIST++ for training. Bailando [32] pioneered using a VQ-VAE for 3D pose encoding, followed by an Actor-Critic GPT model to generate body poses conditioned on music. EDGE [37] applied a diffusion framework to predict human poses from a noisy sequence, conditioning on music and employing long-form sampling for extended dance sequences. Despite these advancements, existing approaches trained on 3D datasets often generate dance poses with limited diversity and are primarily confined to a 3D format. X-Dancer overcomes these data constraints by leveraging a broad spectrum of 2D dance motions, aligning complex movements with musical beats, and utilizing widely available video data to enhance scalability.
2.2 Human Image Animation
With the advancements in diffusion models [14, 12, 4], generating high-quality human videos has become feasible. The introduction of ControlNet [49] and PoseGuider [15] has further empowered prior methods [15, 6, 43, 52] to create realistic dance videos from a reference image using motion signals such as 2D skeletons, dense poses, and depth maps. While these methods provide explicit control over body poses and allow flexible motion manipulation, they often struggle to preserve the original body shape and other identity-specific features, leading to noticeable identity leakage, particularly when significant differences exist between the target and source. Several methods [42, 18, 35] adopt an end-to-end pipeline for generating realistic talking human videos from audio by injecting audio features directly into the network via cross-attention layers. However, while these methods excel in capturing micro-expressions and lip movements, they fall short in handling dynamic, shape-preserving dance movements. X-Dancer uses a diffusion backbone with AdaIN to animate the reference image in an end-to-end framework, enabling high-quality, shape-consistent dance generation while effectively preserving the identity of the reference image.
Figure 2: Overview of X-Dancer. We propose a cross-conditional transformer model to autoregressively generate 2D human poses synchronized with input music, followed by a diffusion model that produces high-fidelity videos from a single reference image $I_{R}.$ First, we develop a multi-part compositional tokenization for 2D poses, encoding and quantizing each body part independently with keypoint confidence. These tokens are then merged into a whole-body, confidence-aware pose using a shared decoder. Next, we train a GPT-based transformer to autoregressively predict future pose tokens with causal attention, conditioned on past poses and aligned music embeddings. For global music style and motion context, we incorporate the entire music sequence and sampled prior poses. With a learnable motion decoder, we generate multi-scale spatial pose guidance upsampled from a learned feature map, incorporating the generated motion tokens within a temporal window (16 frames) using AdaIN. By co-training the motion decoder and temporal modules, our diffusion model is capable of synthesizing temporally smooth and high-fidelity video frames, while maintaining consistent appearance with the reference image with a trained reference net.
3 Method
Given a single portrait as the reference image $I_{R}$ and a conditioning music sequence represented as ${M_{t}},$ our objective is to generate a sequence of dance frames ${I_{M_{t}}}$, where $t=1,…,T$ denotes the frame index. The generated sequence ${I_{M_{t}}}$ seeks to maintain the appearance and background context depicted in $I_{R}$ while present expressive dance movements in harmony with the provided musical rhythm and beats. As illustrated in Fig. 2, our model is trained in two stages: transformer-based 2D human dance motion generation (Section. 3.2), and diffusion-based video synthesis (Section. 3.3) from the generated motion sequence.
3.1 Data Representations
Generative Visual Representations.
To generate human dance videos, we employ Latent Diffusion Models [31] that synthesize samples in the image latent space, facilitated by a pretrained image auto-encoder. During training, latent representations of images are progressively corrupted with Gaussian noise $\epsilon,$ following the Denoising Diffusion Probabilistic Model (DDPM) framework [13, 33]. A UNet-based denoising framework, enhanced with spatial and temporal attention blocks, is then trained to learn the reverse denoising process.
For training, we apply human-centric cropping to the videos, yielding a unified resolution of $896\times 512,$ that encompasses half-body to full-body dances. Instead of directly capturing the intricate pixel-wise movements on music conditions, we first establish the correlation between music and human dance movement, which subsequently guides the final visual synthesis.
Motion Modeling.
In contrast to prior approaches that generate human motions in 3D space, we represent diverse dance motions as 2D pose sequences. Compared to its 3D counterparts, 2D human dance motions are widely accessible from large collections of monocular videos, eliminating the need for complex multiview capture setups or labor-intensive 3D animations. Furthermore, the detection of 2D poses is significantly more reliable and robust. To enhance the realism and motion expressiveness, we model not only large body articulations but also finer details such as head movements and hand gestures, capturing subtle nuances in motion. Notably, we incorporate keypoint detection confidence into our pose representation, allowing the model to account for motion blur and occlusions. Each per-frame whole-body pose with $J$ joints is thus represented as $p\in\mathcal{R}^{J\times 3},$ where the last dimension encodes the keypoint confidence.
Music Embedding.
Inspired from [37], we utilize the pre-trained Jukebox model [9] to extract rich musical features, supplemented by rhythmic information with one-hot encoding of music beats using an audio processing toolbox Librosa [19]. We resample and synchronize these extracted embeddings—denoted as $F^{J}_{1:T}$ for Jukebox and $F^{L}_{1:T}$ for Librosa—to the video frame rate, ensuring per-frame alignment between the music and visual elements.
3.2 Transformer-Based Music-to-Dance Motion Generation
Given a collection of monocular, single-person, music-synchronized dance videos with paired 2D whole-body pose detections, we aim to model the intricate, non-linear correlation between skeletal dance motion and musical features. To achieve this, we first introduce a compositional, confidence-aware VQ-VAE, which captures diverse and nuanced human dance poses across different body parts. Next, we leverage a GPT-like transformer to autoregressively predict future pose tokens, modeling temporal motion transitions within the token space and aligning them with synchronized music embeddings.
Compositional Confidence-Aware Human Pose Tokenization. Our approach builds on the standard VQ-VAE framework, trained in a self-supervised manner. Given whole-body 2D poses with associated keypoint confidences $p$, an 1D convolutional encoder maps $p$ into a latent pose embedding $z_{e}(p)=E(p),$ which is then quantized by mapping to its nearest representation $z_{q}(p)$ within a learnable codebook, and finally the decoder $D$ reconstructs the pose $\hat{p}$ from $z_{q}(p).$ However, as prior studies [32, 45, 41] suggest, the dependencies between spatial keypoints are complex, and a vanilla VQ-VAE often struggles to capture subtle pose details, such as finger movements and head tilts, due to information loss during quantization and multi-frequency nature of pose variations across different body parts.
To improve expressive coverage, we train independent 2D pose encoders $E^{j}$ and learn $B=5$ separate codebooks $Z^{j}_{q}$ for upper and lower half bodies, left and right hands, and head respectively, allowing the model to spatially decompose 2D whole-body pose variations across different frequency levels. By partitioning poses this way, distinct body-part codes can be flexibly combined, enhancing the range of expressiveness represented within each individual codebook. To capture part-wise spatial correlations and ensure information flow across body parts, we concatenate the quantized pose latents and feed them into a shared decoder. The resulting embedding is then mapped to reconstructed keypoint coordinates through separate projection heads, enabling joint reconstruction while preserving nuanced part-specific details.
We train encoder and decoder simultaneously with the compositional codebooks with the following loss function:
$\mathcal{L}_{\text{VQ}}=\sum_{j=1}^{B}\|\hat{p}^{j}-p^{j}||_{2}+\beta\sum_{j=1}^{B}\|\mathrm{sg}\left[z^{j}_{e}(p)\right]-z^{j}_{q}(p)\|_{2},\vspace{-2mm}$ (1)
where $\mathrm{sg}$ is a stop gradient operation, and the second term is a commitment loss with a trade-off $\beta.$ We utilize exponential moving average (EMA) [40] to update codebook vectors and remove the embedding loss $\sum_{j=1}^{B}\|z^{j}_{e}(p)-\mathrm{sg}\left[z^{j}_{q}(p)\right]\|_{2}$ for more stable training.
Cross-Conditional Autoregressive Motion Modeling. After training the compositional quantized multi-part codebooks, 2D human poses detected in our training videos can be represented as sequences of codebook tokens via encoding and quantization. Specifically, each detected pose is mapped to a sequence of corresponding codebook indices, structured as one-hot vectors indicating the nearest codebook entry for each element. We denote this as $C^{j}_{1:T}=((c^{j}_{1,1},…,c^{j}_{K,1}),\ldots,(c^{j}_{1,T},\ldots,c^{j}_{K,T}))$, where $K$ is the number of tokens per body part $j$ in each frame.
With this quantized motion representation, we design a temporal autoregressive model to predict the multinomial distributions of possible subsequent poses for each body part, conditioned on both the input music embeddings $F^{J}_{1:T}$ and $F^{L}_{1:T}$ and the preceding motion tokens. Our motion generation transformer is conditioned on the music inputs in two ways. First, we combine the Jukebox and Librosa music embeddings to form starting tokens, which serves as a global music context, such as styles and genres, that informs the entire motion sequence generation. Second, inspired by [32], we concatenate the frame-wise projected music embeddings with the corresponding motion tokens as inputs to the transformer model, ensuring precise synchronization between the motion and music features. This dual conditioning allows our model to produce both globally coherent and locally synchronized dance movements. We use $T=64$ frames for our autoregressive model training. To handle extended motion sequence generation with consistent motion styles and smooth transitions, we additionally incorporate cross-segment motion context into the transformer model. Specifically, we uniformly sample a subset of 8 frames from the previous motion segment as the motion context, and append them after the global music context inputs. We denote the combined context as $F_{g}.$
Since we model body parts independently, maintaining coherence in the assembled whole-body poses is essential to avoid asynchronous movements (e.g., upper and lower body moving in different directions). To address this, we leverage mutual information across multi-part motions, designing our model with cross-conditioning across body parts. Specifically, we employ a GPT model to estimate the joint distributions of $C^{j}_{1:T}$ as follows,
$\phi(C^{1:B}_{1:T}|F_{g})=\prod_{t=1}^{T}\prod_{j=1}^{B}\prod_{k=1}^{K}$
$\phi\bigl{(}c^{j}_{k,t}|c^{1:B}_{1:K,<t},c^{<j}_{1:K,t},c^{j}_{<k,t},F^{J}_{\leq t},F^{L}_{\leq t},F_{g}\bigr{)}\vspace{-2mm}$ (2)
We structure the cross-conditioning between body parts in two ways: (1) the current motion token is conditioned on all preceding motion information from all body parts, ensuring inter-part temporal coherence; (2) by ordering body parts as upper and lower body, head, and hands, we build the hierarchical dependencies from primary components (upper/lower body) to finer, high-frequency movements (head and hands). Since each body part’s pose is represented with a small set of tokens, we empirically observe that causal attention is sufficient to model the next-token distribution within each part. This modeling strategy preserves motion coherence of each body part as a whole, producing expressive and plausible dance movements.
Our GPT model is optimized through supervised training using a cross-entropy loss on the next-token probabilities. Notably, because our pose tokenization includes associated keypoint confidences, the transformer also learns to model temporal variations in confidence, such as those caused by motion blur and occlusions, enabling it to capture more realistic motion distributions observed in videos.
Figure 3: Qualitative Comparisons. Among all the methods, X-Dancer achieves the most expressive and high-fidelity human dance video synthesis, maintaining the highest consistency with both the reference human characteristics and the background scene.
3.3 Diffusion-Based Dance Video Synthesis
With the generated motion token sequences from our trained transformer model, we employ a diffusion model to synthesize high-resolution, realistic human dance videos, conditioned on the generated motions and a given reference image. To achieve this, we leverage a pretrained T2I diffusion backbone [31] and incorporate additional temporal modules [12] to improve cross-frame consistency and ensure temporal smoothness. For transferring the reference image context, a reference network, as a trainable copy of the backbone UNet, extracts reference features of identity appearance and background which are cross-queried by the self-attentions within the backbone UNet. Motion control is achieved through an addition module, often configured in ControlNet [49] or light-weight PoseGuider [15], which translates the motion conditions into 2D spatial guidance additive to the UNet features.
To incorporate the generated motion tokens into human image animation, one approach is to utilize the trained VQVAE decoder $D$ to decode pose tokens into keypoint coordinates, which are then visualized as 2D skeleton maps for motion conditioning. These skeleton maps can provide motion guidance to the final diffusion synthesis through a PoseGuider module. While effective to some extent, this approach introduces a non-differentiable process due to the skeleton map visualization, which impedes end-to-end training and often results in the loss of keypoint confidence information. Additionally, since the temporal module is trained on real video data where pose sequences are typically smooth, it may struggle with jittery or inconsistent motions generated by the transformer model at inference.
In place of explicit translation of motion tokens into 2D skeleton maps as pose guider conditions, we introduce a trainable motion decoder that implicitly and directly translates the 1D pose tokens into 2D spatial guidance, added towards the intermediate features of the denoising UNet. Starting from a learnable 2D feature map, the decoder injects the 1D pose token sequences including the keypoint confidences with AdaIN layers [16], progressively upsampling this feature map into multiple scales aligned with the resolutions of the denoising UNet features. The motion decoder is trained alongside the temporal module within a 16-frame sequence window, effectively decoding token sequences into continuous pose guidance that integrates temporal context from adjacent frames. Moreover, by incorporating reference image context during training, we observe empirically that the synthesized pose guidance retains minimal identity and body shape information compared to skeleton-based pose guiders, enabling generated pose tokens to adapt seamlessly to characters with varied body shapes and appearances.
4 Experiments
Table 1: Quantitative comparison on motion generation.
Metrics AIST++ Dataset / In-House Dataset
FVD $\downarrow$ DIV $\uparrow$ BAS $\uparrow$
Ground Truth 509.58/129.75 34.10/29.67 0.24/0.22
Hallo [42] 548.81/249.12 28.66/28.98 0.16/0.20
Bailando [32] 621.22/534.02 22.34/24.05 0.19/0.19
EDGE [37] 639.46/303.36 24.87/27.29 0.26/0.24
X-Dancer 531.52/238.22 25.61/28.08 0.23/0.21
Table 2: Quantitative comparison on visual synthesis.
Metrics In-House Dataset
FVD $\downarrow$ FID-VID $\downarrow$ ID-SIM $\uparrow$
Hallo [42] 609.08 76.99 0.4870
Bailando [32] + PG 583.26 100.02 0.3392
EDGE [37] + PG 613.81 93.73 0.3034
Our motion + PG 735.05 72.71 0.4894
X-Dancer 507.06 61.94 0.5317
4.1 Implementation Details
Dataset. Our model is trained on a curated visual-audio dataset comprising 107,546 monocular recordings of human dance performances in both indoor and outdoor settings, with an average clip duration of 30 seconds. Each video is cropped to a resolution of $896\times 512$ and resampled to $30$ fps, encompassing a broad spectrum of half- to full-body dances across diverse music genres and showcasing individuals with varied body shapes and appearances.
Training. We train our full pipeline in three stages. First, we train a multi-part pose VQ-VAE to encode and quantize 60 joint coordinates, with keypoint confidence scores, into 5-part pose tokens. The pose of each body part uses 6 tokens, each with a unique 512-entry codebook of 6-dimensional embeddings. This VQ-VAE is trained for 40k steps with a batch size of 2048 at a learning rate of $2\times 10^{-4}$. Next, we train an autoregressive model for pose token prediction with 300k steps, initialized with pretrained GPT-2 model weights, using a batch size of 24 and a learning rate of $1\times 10^{-4}$. The model is trained with 64-frame pose sequences with a context window of 2224 tokens in total, enabling extended dance motion generation via sliding segments during inference. In the final diffusion stage, we fine-tune the denoising UNet and ReferenceNet with two randomly selected frames from a video, followed by co-training the motion guidance decoder and the temporal module with diffusion losses on consecutive 16 frames. We train the diffusion stage with 90k steps at a learning rate of $1\times 10^{-5}$ and a batch size of 16. All stages are trained with 8 A100 GPUs using the AdamW optimizer [44].
Inference. For inference, we initiate our autoregressive dance motion generation from pose tokens encoded from the reference image pose, maintaining skeleton consistency with the specified body shape. Extended dance sequences are generated in 64-frame sliding segments with a 12-frame overlap, while 8 uniformly sampled frames from previous generated motions serve as global motion context. Using the full generated pose token sequences, we synthesize all video frames simultaneously, applying prompt traveling [36] to improve temporal smoothness.
Table 3: Ablation on our transformer model designs.
FVD $\downarrow$ DIV $\uparrow$ BAS $\uparrow$
w/o global music context 265.73 27.04 0.2142
w/o global motion context 247.54 26.42 0.2154
sub-dataset + GPT-medium 402.63 24.40 0.2112
sub-dataset + GPT-small 332.93 24.58 0.2046
X-Dancer 238.22 28.08 0.2182
4.2 Evaluations and Comparisons
To the best of our knowledge, no existing approach addresses music-driven 2D video generation from a single human image. For baseline comparisons, we adapted and combined established models to create two baselines. First, we adapted the audio-driven, diffusion-based portrait animation model Hallo [42], retraining it for our task by substituting its audio features with our music embeddings to animate human images via cross-attention modules. For the second baseline, we utilize 3D music-to-dance generation models like Bailando [32] and EDGE [37] for motion synthesis, projecting their outputs into 2D skeleton maps, which are then fed into a diffusion model with a pose guider for controlled image animation. We note that during our submission, the AIST dataset [38] was unavailable due to maintenance, and we were unable to train our model on the same dataset as Bailando [32] and EDGE [37]. We conduct separate evaluations of all models on the test set of AIST++ [21] (40 videos) as well as on a test split of our curated music-video dataset (230 videos).
Quantitative Evaluation. We numerically compare our method with baselines in terms of quality of both motion generation and video synthesis. Specifically, we calculate the Fréchet Video Distance (FVD) [39, 10] between generated dance motions and ground-truth training sequences for assessment of motion fidelity. For motion diversity, we compute the distributional spread of generated pose features (DIV) [21, 32], for which the generated motion should aim to match the scores of the ground truth distribution rather than maximize their absolute values. To numerically evaluate the alignment between the music and generated dancing poses, we follow [21, 32] to measure the Beat Align Score (BAS) by the average temporal distance between each music beat and its closest dancing beat. These evaluations are conducted in 2D pose space, where we project the 3D skeleton motion of Bailando and EDGE into 2D and detect poses over synthesized videos from the retrained Hallo model. As shown in Tab. 2, our method surpasses all baselines in terms of motion fidelity (FVD), while achieving the second-best music beat alignment (BAS) and diversity (DIV). Notably, EDGE is trained on sequences from professional dancers, whereas our dataset comprises videos of everyday individuals, reflected in the beat alignment score of the ground-truth videos. Despite this, our method significantly outperforms Bailando trained on AIST++ and Hallo trained on our curated dataset for beat alignment. Hallo entangles motion generation and video synthesis, and achieves a higher DIV score than X-Dancer primarily due to its extremely noisy video outputs which results in jittering and chaotic skeleton motions following pose projection.
Figure 4: Human image animation after finetuning motion transformer with 30 dance videos of Subject Three.
For evaluation of video synthesis fidelity, we measure the FVD and FID-VID [3] score between the ground-truth and generated dance videos. Additionally, we assess identity preservation using the ArcFace score [8], which measures the cosine similarity of face identity features (ID-SIM). All metrics are evaluated on our test video dataset. As an extra baseline, we replace the motion generator in EDGE and Bailando with our trained transformer model and generate the animation using a pose guider. As shown in Tab. 2, our method achieves the highest visual quality and identity preservation, which we attribute to our disentangled design of motion generation and video synthesis (compared to Hallo) and the use of an implicit motion token decoder rather than an explicit pose guider.
Qualitative Comparisons. We present visual comparisons between our method and the baselines in Fig. 3. For more dynamic comparisons and user study on perceptual quality, please refer to our supplementary material. The modified Hallo [42] represents an end-to-end diffusion pipeline that directly synthesizes the final video without intermediate motion generation. However, it exhibits noticeable artifacts, particularly in large articulated body motions, and often fails to preserve the human body’s articulation topology. Bilando [32] and EDGE [37] generate body motion in 3D space without considering the scene context or the human shape in the reference image. Despite post-processing for skeleton alignment in the pose guider input, these methods still struggle with significant identity and shape distortions, often producing unnatural interactions with the background scene. Furthermore, they do not model head and hand motions, leading to more rigid and less expressive dance movements compared to X-Dancer.
Finetuning for Characterized Choreography. While our method operates as a zero-shot pipeline, generalizing seamlessly to new reference images and music inputs, it can also be fine-tuned for characterized choreography using only a few sample dance videos. This adaptability is challenging for 3D motion generation models like EDGE [37] and Bailando [32], which require extensive effort in creating 3D dance movements. As shown in Fig. 4 and our supplementary video, our method successfully captures and mimics the specific choreography after fine-tuning with only 30 dance videos from diverse performers, showcasing its efficiency and versatility in adapting to specific dance styles.
4.3 Ablation Study
We conduct ablation studies by systematically removing individual design components from our full training pipeline. First, we assess the effectiveness of the multi-part VQ-VAE in capturing subtle and expressive human movements. Compared to a single-part whole-body VQ-VAE, our approach achieves a significantly lower L1 reconstruction error per joint (0.5px vs. 0.83 px), with pronounced improvements in the hands (0.4px vs. 0.64px) and head (0.42px vs. 0.52px). This improvement translates into superior motion diversity, expressiveness, and control precision in the diffusion model, as shown through quantitative metrics and visual examples in our supplementary material. Next, we assess the impact of global music and motion context on motion generation. As presented in Tab. 3, both contexts contribute to producing more consistent, plausible, and music-synchronized motions. We further analyze the benefits of 2D motion modeling and transformer-based autoregressive generation by scaling both model parameters and dataset size. Across all metrics, we observe significant performance gains (Tab. 3) as the number of training parameters increases from 117M (GPT-small) to 345M (GPT-medium) and data scale from 10k to 100k videos, underscoring the scalability potential of monocular dance video data with our architecture, sheding light on further performance improvements as they scale. In Tab. 2, we compare our token-based motion guidance with AdaIN against a pose guider with explicitly decoded skeleton map. Our method demonstrates reduced motion jittering and enhanced temporal consistency due to the temporal motion context, and showcases superior identity and body shape preservation, as also visually shown in our supplementary video.
5 Conclusion
We present X-Dancer, a novel framework that unites an autoregressive transformer with a diffusion model to generate high-quality, music-driven human dance videos from a single reference image. Unlike prior works, X-Dancer models and generates dance movements in 2D space, harnessing widely accessible 2D poses from monocular dance videos to capture diverse, expressive whole-body motions. Our method achieves state-of-the-art results in video quality, motion diversity and expressiveness, providing a scalable and adaptable solution for creating vivid, music-aligned dance videos across various human forms and styles.
Limitations and Future Work. Our model is trained solely on curated real-human daily dance videos from internet, which can be noisy and the dance movements may not exhibit the precision found in professional dancer videos. Consequently, out-of-domain human images may lead to rendering artifacts, and the generated dance motions may occasionally lack alignment with music. While we designed our pipeline to be end-to-end trainable and scalable, we currently implement it in stages due to memory limitations. Future work will explore large-scale, multi-machine training to further enhance performance and efficiency.
Ethics Statement. Our work aims to improve human image animation from a technical perspective and is not intended for malicious use like fake videos. Therefore, synthesized videos should clearly indicate their artificial nature.